OpenStudy (anonymous):

3x - 4y = 7 3x + 2y = -5 When the second equation is subtracted from the first, the resulting equation is _____. Be sure to show all work!

OpenStudy (anonymous):

3x - 3x = 0 -4y - 2y = -6y 7 - -5 = 12 and we get -6y = 12

OpenStudy (anonymous):

good job